If given:
A=68 a= ?
B= ? b= 8
C= ? c=5
Determine C
A.) +31.38+
B.) +32.92+
C.) +33.12+
D.) +37.02+

Found 2 solutions by Boreal, KMST:
Answer by Boreal(15235)

a^2=b^2+c^2-2bc cos A
=64+25-2(40) *0.3746=89-29.97=59.03
sqrt(59.03)=7.68
Now use law sines
sin 68/7.68=sinC/5
sin C=.6034
Take arc sin both sides: C=37.11
I will choose D.

Answer by KMST(5328)

We are seeing a second wave of problem featuring this triangle.
Similar questions asked previously (and maybe the same ones)
would be easier to find if placed under the headings of
"Triangles" (Geometry) or
"Trigonometry" (Algebra II).
I searched here and found a few recent ones,
and I knew there was an earlier wave.
I even found an earlier posting of this problem
with the answer "duplicate."
They were all under "test."

I knew I had seen this problem answered here before,
so I searched online for
"If given: A=68 a= ? B= ? b= 8 C= ? c=5"
I found:
this problem with its solution,
a previous posting of the same problem with an answer, and
a PDF file from Houston Independent School District
(PreCalc_Unit_6_Soln.pdf)
that uses that same triangle in
example 15 (first just the problem, followed by some instructions, and then solution values for C, B, and a on the next page).